[1/6 + 1/2] / [19/42 / 1/6] = what
Answer by Maths68(1474)

=[1/6 + 1/2] / [19/42 / 1/6]
(Take LCM of Numerator and Change division to multiplication of Denominator)
=[[(1*1)+(1*3)]/6] / [(19/42) * (6/1)]
=[(1+3)/6] / [(19/42) * (6/1)]
=(4/6) /(19/7)
=(2/3)/(19/7)
Again Chang Division to multiplication so denominator will be inverse)
=(2/3)*(7/19)
=(2*7)/(3*19)
=14/57 Ans.

or another way to do this
let
a= 1/6 + 1/2
b= (19/42) / (1/6)
Solve 'a' first
a= 1/6 + 1/2
Take LCM
=[(1*1)+(1*3)]/6
=(1+3)/6
=4/6
=2/3
Now Solve 'b'
b= (19/42) / (1/6)
Change Division into Multiplication then Denominator will be inverse
b= (19/42) * (6/1)
= (19*6)/42
= 114 / 42
=19/7
Now we have
= a/b
Put the values of a and b
= (2/3)/(19/7)
Again change division to Multiplication then denominator will be inverse
= 2/3 * 7/19
= 14/57 Ans.